OpenStudy (anonymous):

what is the value of f(0) for the function...

OpenStudy (anonymous):

\[\log_{10} 10+9^{x}+(x-2)(x-1)\]

OpenStudy (amoodarya):

do you mean ? \[f(x)=\log_{10} (10+9^x+(x-1)(x-2))\]

OpenStudy (anonymous):

hmm... no thats not how it asked me but could it have been meant that way?

OpenStudy (phi):

you answer by replace x with 0 in your expression

OpenStudy (phi):

\[ \log_{10} 10+9^{0}+(0-2)(0-1) \] let's do the easy part: what is 0-2 ? what is 0-1 ?

OpenStudy (anonymous):

\[\log_{10} 10+9^0(-2)(-1)\]

OpenStudy (phi):

I hope you mean \[ \log_{10} 10+9^{0}+(0-2)(0-1)\\ \log_{10} 10+9^{0}+-2\cdot -1 \]

OpenStudy (anonymous):

yes

OpenStudy (phi):

next, what is -2*-1 ?

OpenStudy (anonymous):

2

OpenStudy (phi):

so we now have \[ \log_{10} 10+9^{0}+ 2 \] next: any number (except 0) raised to the "0 power" is 1. using that rule, what is 9^0 ?

OpenStudy (anonymous):

1

OpenStudy (phi):

so \[ \log_{10} 10+1+2 \\ \log_{10} 10+3 \]

OpenStudy (phi):

I assume you don't know the log base ten of 10 ? log means "what exponent" base ten to "what exponent" is 10 ? 10^x = 10 what is x ?

OpenStudy (phi):

you should know 10^0 = 1 10^1 = 10 10^2 = 100

OpenStudy (anonymous):

so \[\log_{10} 10\] equals to 1?

OpenStudy (phi):

yes. If you have a calculator and try log 10 you should get 1 so we now have 1+3

OpenStudy (anonymous):

ok got it